# Phase diagram of two-component mean-field Bose mixtures

### v1

@@ V Liquid-gas type transition within the normal phase
As reported in the recent study of Ref. [32], for sufficiently large, positive values of $a_{12}$, the two-component Bose mixture may undergo a phase transition which does not involve condensation, but manifests itself in a jump of concentrations of the mixture constituents (both remaining below the critical condensation density). Interestingly, a somewhat similar transition was identified [30] via Monte Carlo simulations of a two-dimensional Bose mixture in a completely different parameter regime characterized by attractive interspecies couplings. Below, we demonstrate that within the present mean-field model in dimensionality $d=3$, the occurrence of this transition is completely excluded as long as $D>0$. We identify the parameter ranges (corresponding to sufficiently large positive $a_{12}$ or $T$), where it does occur in the case $D<0$. We also find that there exists a range of system parameters where the two normal phases coexist with the two BEC phases, implying the presence of quadruple points.
Anticipating the possibility of first-order transitions terminating at critical points, and occurring within the non-condensed region of the phase diagram, we employ a method analogous to the approach of the previous section with the aim of locating the critical points. We express $n_{1}$ as a function of $n_{2}$ using Eq. (12), which enables us to define $\tilde{\Phi}(n_{2})$. At the critical point, the second derivative of $\tilde{\Phi}(n_{2})$ must vanish. Unlike the case analyzed in the previous section, we do not have an explicit formula for the coexistence line. In the general situation, the missing equation is provided by the condition of vanishing third derivative of $\tilde{\Phi}(n_{2})$. The set of equations to tackle the problem involves two equations corresponding to the saddle point conditions and two conditions involving higher derivatives: $\tilde{\Phi}^{\prime\prime}(n_{2})=\tilde{\Phi}^{\prime\prime\prime}(n_{2})=0$. The condition $\tilde{\Phi}^{\prime\prime}(n_{2})=0$ can be cast in the following form:
@@
Figure 4: Evolution of the tricritical points and the critical point upon varying temperature at fixed $\frac{a_{12}}{a}=5$ ($\lambda_{1}=\lambda_{2}=\lambda$, $a_{1}=a_{2}=a$). The subsequent illustrations correspond to $\beta a\lambda^{-3}=0.02$, $\beta a\lambda^{-3}\approx 0.039$, $\;\beta a\lambda^{-3}\approx 0.044$, and $\beta a\lambda^{-3}=0.09$. The orange discs represent the (putative) tricritical points, the orange triangles mark the (putative) critical points, and the black discs indicate the triple points. The orange points are physical if they are located to the left (and below) of the triple point.
@@ VI.1 $D<0$
$\mu_{1}-\frac{a_{12}}{a_{2}}\mu_{2}<0.$ (30)
If $\frac{a_{2}}{a_{12}}\mu_{1}>\mu_{2}$, the BEC1 phase occurs alongside BEC2. In the complementary case, we find that when $\mu_{2}>\frac{a_{12}}{a_{1}}\mu_{1}$, type-2 particles condense along type-1 particles.

### v2

@@ V Liquid-gas type transition within the normal phase
As reported in the recent study of Ref. [34], for sufficiently large, positive values of $a_{12}$, the two-component Bose mixture may undergo a phase transition which does not involve condensation, but manifests itself in a jump of concentrations of the mixture constituents (both remaining below the critical condensation density). There is no fundamental difference between the two non-condensed phases. This transition, characterized by a jump in densities, is analogous to the liquid-gas transitions occurring in classical fluids. Interestingly, a somewhat similar transition was identified [63] via Monte Carlo simulations of a two-dimensional Bose mixture in a completely different parameter regime characterized by attractive interspecies couplings. Below, we demonstrate that within the present mean-field model in dimensionality $d=3$, the occurrence of this transition is completely excluded as long as $D>0$. We identify the parameter ranges (corresponding to sufficiently large positive $a_{12}$ or $T$), where it does occur in the case $D<0$. We also find that there exists a range of system parameters where the two normal phases coexist with the two BEC phases, implying the presence of quadruple points.
Anticipating the possibility of first-order transitions terminating at critical points, and occurring within the non-condensed region of the phase diagram, we employ a method analogous to the approach of the previous section with the aim of locating the critical points. We express $n_{1}$ as a function of $n_{2}$ using Eq. (12), which enables us to define $\tilde{\Phi}(n_{2})$. At the critical point, the second derivative of $\tilde{\Phi}(n_{2})$ must vanish. Unlike the case analyzed in the previous section, we do not have an explicit formula for the coexistence line. In the general situation, the missing equation is provided by the condition of vanishing third derivative of $\tilde{\Phi}(n_{2})$. The set of equations to tackle the problem involves two equations corresponding to the saddle point conditions and two conditions involving higher derivatives: $\tilde{\Phi}^{\prime\prime}(n_{2})=\tilde{\Phi}^{\prime\prime\prime}(n_{2})=0$. By elementary algebraic manipulations, the condition $\tilde{\Phi}^{\prime\prime}(n_{2})=0$ may be cast in the following form:
@@
Figure 4: Evolution of the tricritical points and the critical point upon varying temperature at fixed $\frac{a_{12}}{a}=5$ ($\lambda_{1}=\lambda_{2}=\lambda$, $a_{1}=a_{2}=a$). The subsequent illustrations correspond to $\beta a\lambda^{-3}=0.22$, $\beta a\lambda^{-3}\approx 0.16$, $\beta a\lambda^{-3}\approx 0.15$, and $\beta a\lambda^{-3}=0.11$. The orange discs represent the (putative) tricritical points, the orange triangles mark the (putative) critical points, and the black discs indicate the triple points. The orange points are physical if they are located to the left (and below) of the triple point.
@@ VI.1 $D<0$
$\mu_{1}-\frac{a_{12}}{a_{2}}\mu_{2}<0,$ (30)
Here, we have used the relation $D=a_{1}a_{2}-a_{12}^{2}$ along with the fact that in aforementioned limit, $n_{2}\sim\frac{\mu_{2}}{a_{2}}$. If $\frac{a_{2}}{a_{12}}\mu_{1}>\mu_{2}$, the BEC1 phase occurs alongside BEC2. In the complementary case, we find that when $\mu_{2}>\frac{a_{12}}{a_{1}}\mu_{1}$, type-2 particles condense along type-1 particles.